# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

### v1

@@ V-B Results on UniAttackData+
We employed a range of CLIP-based baseline models (including CLIP-V, CLIP, CoOp without class-specific context, and CoOp with class-specific context) to evaluate the performance of shared-space learning models on our dataset. In addition, we assessed algorithms specifically designed for PAD/FFD, such as CFPL-FAS, MoE-FFD, and LLA-Net, to determine their generalization capabilities. We also compared our approach with the recently released $FA^{3}-CLIP$, an algorithm specifically designed for unified attack detection. The detailed experimental results are presented in Table VII.
Authenticity-Only Testing Protocol $p1$:
This protocol primarily aims to evaluate the model’s true capability in distinguishing between real and fake samples. As shown in the experimental results table (Table VII), models that attempt to update the encoder generally achieve better performance. This is because, within UniAttackData+, a shared feature space still exists where different attack types can be clustered, and enhancing the image encoder’s feature extraction capability improves model performance. However, these models have a large number of parameters, making them unsuitable for deployment in real-world environments.
Conversely, CoOp, which relies on prompt matching, attempts to map fixed features to relevant regions. However, given the vast and unpredictable variety of attack types, a single prompt is insufficient to address the diversity and sophistication of these attacks. Models specifically designed for PAD and FFD focus solely on the manipulated regions of the respective images, neglecting the common forgery cues and underlying processes shared across UAD tasks. Notably, CDCN++, a purely convolutional model, also demonstrates decent performance, highlighting the importance of local region analysis in UAD tasks. This suggests that local feature extraction remains crucial for robust attack detection.
Our HiPTune adopts a visual prompt approach, which not only guides the encoder in feature extraction but also maintains model efficiency. Furthermore, our hierarchical prompt tree encapsulates multi-level, fine-grained classification criteria, achieving highly precise classification performance.
Real-World Robustness Testing Protocol $p2$:
This protocol is widely applicable in real-world scenarios, aiming to evaluate the model’s performance on seen attack types (such as physical 2D attacks) and unseen attack methods (such as face-swapping techniques like Roop and SimSwap). Similar to protocol p1, models that train the encoder generally achieve superior performance. This is because identifying a shared feature space helps the model generate a general mapping for previously unseen methods, effectively leveraging existing knowledge to handle new attack techniques.
In contrast, models like CoOp, which rely on prompt matching, are more prone to overfitting the training data. As a result, their detection performance significantly declines when encountering new data distributions. For instance, CDCN++, a model specifically designed for PAD, utilizes a central difference convolutional structure that emphasizes local detail information, achieving the best results in our benchmark experiments. This finding highlights the importance of focusing on local detail information to ensure model generalization in UAD tasks.
Our HiPTune model adopts a visual prompt approach, introducing prompts on the image encoder side to guide the frozen encoder in extracting relevant features while maintaining efficiency. Unlike conventional prompt-matching methods, our prompts serve only as guidance, avoiding excessive dependence on existing knowledge for decision-making. This design not only enhances the model’s generalization capability against unseen attack methods but also ensures robustness and interpretability.
Upward Compatibility Protocol $p3.1$:
This protocol aims to assess whether models trained on simple attack types can maintain high robustness in testing environments featuring complex and advanced attacks. We observe a sharp performance decline in both CoOp models, which we attribute to the absence of encoder guidance. Relying solely on text-side mapping relationships fails to achieve strong reasoning and generalization capabilities.
In contrast, our HiPTune model, which adaptively integrates different types of prompts, demonstrates superior performance. Despite being trained on limited samples, HiPTune can make judgments based on existing higher-level prompts while leveraging data from other attack types to assist in inference. This design ensures that the model retains strong generalization even in the presence of advanced attack types.
Downward Compatibility Protocol $p3.2$:
This protocol aims to evaluate whether models can leverage advanced knowledge to accurately detect simpler attack types. As shown in the table, most models exhibit poor performance, indicating that existing methods are insufficient to effectively utilize anti-forgery knowledge derived from advanced digital generative attacks. This also reveals the potential non-transferability between these two types of knowledge.
In contrast, our HiPTune model demonstrates superior performance. It effectively borrows knowledge from higher-level prompts and dynamically combines existing classification criteria, highlighting its advanced design. This adaptive approach allows HiPTune to generalize knowledge from complex attacks to simpler ones, ensuring robust performance across diverse scenarios.
V-C Results on UAD-Benchmark
JFSFDB-UniAttackData+ UniAttackData-UniAttackData+ JFSFDB+UniAttackData-UniAttacData+
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 35.83 67.91 53.44 35.83 41.41 60.9 49.69 41.41 39.81 63.4 54.1 39.81
CLIP 39.32 63.95 55.65 39.32 40.58 62.75 50.38 40.58 39.36 64.08 51.92 39.36
CoOp(unified) 39.59 63.56 58.29 39.6 26.93 78.77 70.44 26.93 29.77 76.77 67.4 29.77
CoOp(specific) 39.39 63.89 58.71 39.39 28.62 77.5 69.9 28.62 30.3 75.9 68.49 30.3
HiPTune 32.88 69.34 60.51 32.88 20.16 88.53 86.82 20.23 25.93 83.05 73.82 25.94
UniAttackData-JFSFDB UniAttackData+-JFSFDB UniAttackData+UniAttacData+-JFSFDB
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 36.16 64.97 70.81 36.16 51.65 48.75 47.88 51.65 45.12 56.26 50.52 45.12
CLIP 36.98 64.05 76.04 36.98 50.72 48.25 44.09 50.72 44.42 58.1 54.11 44.42
CoOp(unified) 40.02 64.12 58.53 40.02 43.55 59.51 54.28 43.55 38.21 65.84 59.44 38.21
CoOp(specific) 39.44 64.83 59.02 39.44 42.95 59.99 55.39 42.95 38.13 66.04 60.67 38.13
HiPTune 34.39 71.23 80.31 34.42 28.14 79.22 70.46 28.78 32.93 70.89 72.23 32.92
JFSFDB-UniAttackData UniAttackData+-UniAttackData JFSFDB+UniAttackData+-UniAttacData
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 29.94 75.58 62.11 29.94 22.26 84.22 56.92 22.26 13.15 91.86 77.74 13.15
CLIP 29.27 80.12 67.63 29.27 22.38 84.11 75.76 22.38 16.59 90.14 82.5 16.59
CoOp(unified) 37.17 67.2 58.91 37.17 25.4 80.5 71.77 25.4 30.28 76.11 65.41 30.28
CoOp(specific) 41.4 61.95 57.39 41.4 26.37 79.32 70.68 26.37 27.34 79.19 67.3 27.34
HiPTune 21.36 85.59 74.39 31.56 18.62 90.75 88.64 18.63 8.55 94.67 90.26 8.56
TABLE VIII: Results on existing UAD datasets.
In the UAD-Benchmark, training and testing models across different datasets allow us to evaluate not only their generalization to various attack types but also their domain generalization capabilities. By examining the table, we can draw two key insights:
(1) As shown in the second column of the table, models trained on UniAttackData+ demonstrate strong anti-forgery capabilities that can be effectively transferred to other datasets. This highlights the robustness and generalizability of the knowledge learned from UniAttackData+.
(2) Baseline models show limited improvement, suggesting that we should design models specifically tailored to the characteristics of the dataset. Such models should be capable of fully leveraging the knowledge contained within the dataset, maximizing its utility for model training and evaluation.
V-D Results on JFSFDB
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 95.73 100 68.87 85.33 44.54 68.05 65.54 54.72 65.1 71.99
Xception 99.88 100 94.47 90.62 63.88 90.15 98.21 64.86 84.27 87.37
MultiAtten 99.56 100 68.68 89.21 36.88 72.66 70.24 63.37 64.34 73.88
CDCN++ 100 99.56 94.51 82.65 47.66 76.76 97.53 56.65 78.34 81.51
DeepPixel 100 100 89.4 95.17 56.87 80.6 99.26 67.54 85.51 86.04
ResNet50 99.99 100 93.71 94.33 58.95 88.79 97.04 62.35 82.51 86.41
ViT 100 100 95.45 95.13 35.6 89.12 98.65 73.79 86.61 86.04
CLIP-V 99.99 100 96.85 97.29 76.62 93.6 93.05 82.09 75.4 90.54
CLIP 99.99 100 95.58 95.82 80.4 92.89 95.92 86.08 79.4 91.79
CoOp(Unified) 99.86 100 94.42 83.73 91.94 94.63 79.49 64.46 62.05 85.62
CoOp(Specific) 99.83 100 93.72 84.14 92.61 92.89 81.41 66.63 61.88 85.9
Sep. HipTune 99.97 100 98.62 98.62 95.87 98/62 100 89.72 89.63 93.12
MesoNet 92.55 97.94 78.42 84.01 45.87 70 56.66 50.57 55.03 72.56
Xception 99.8 100 97.62 91.56 75.49 79.38 96.69 63.69 78.87 87.01
MultiAtten 99.38 100 78.34 92.89 58.31 77.51 69 58.11 60 77.06
CDCN++ 99.99 98.89 96.8 78.93 61.54 76.39 95.79 56.42 76.92 82.41
DeepPixel 100 100 97.31 93.97 66.65 74.58 99.22 66.36 88.45 87.39
ResNet50 99.93 100 98.1 93.5 79.64 86.03 98.83 67.18 85.55 89.86
ViT 100 100 100 95.85 62.74 91.88 98.83 75.11 88.36 90.31
JFSFDB-Uni 100 100 99.96 97.99 67.1 95.57 99.12 75.9 89.76 91.71
CLIP-V 100 100 95.91 97.23 71.26 83.96 95.69 82.84 79.64 89.44
CLIP 100 100 97.48 96.39 86.13 81.07 94.83 84.06 78.99 90.83
CoOp(Unified) 99.71 100 95.82 81.58 89.27 93.62 78.55 68.56 64.88 85.78
CoOp(Specific) 99.78 100 95.33 80.68 92.82 93.34 76.42 67.58 65.06 85.67
AUC(%) Joint HipTune 100 100 100 99.62 95.43 97.82 100 90.88 93.35 97.46
TABLE IX: Results on JFSFDB AUC.
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 11.67 0 35.12 21.9 55.82 36.81 38.93 46.16 38.78 31.69
Xception 1.14 0 11.31 18.57 42.3 18.13 5.54 38.92 24.07 17.78
MultiAtten 2.78 0 35.71 18.57 61.22 33.82 35.54 39.73 40.46 29.76
CDCN++ 0.14 3.33 11.9 22.8 52.76 32.01 7.5 44.75 28.19 22.6
DeepPixel 0 0 17.86 15.71 46.42 28.23 2.32 36.82 22.81 18.91
ResNet50 0.36 0 13.69 12.38 45.24 19.94 8.18 40.61 26.01 18.49
ViT 0.14 0 11.31 11.9 60.4 19.47 4.64 32.75 22.76 18.15
CLIP-V 0.49 0 7.74 8.58 29.56 13.17 13.66 25.43 29.89 14.28
CLIP 0.34 0 11.51 12.86 26.74 14.61 11.47 23.14 27.16 14.2
CoOp(Unified) 1.51 0 12.7 25.95 17.02 12.06 27.14 39.42 41.21 19.67
CoOp(Specific) 1.66 0 15.08 25.48 15.68 15.25 25.98 37.69 41.53 19.82
Sep. HipTune 0.35 0 4.58 6.56 11.57 12.82 1.89 20.86 28.77 9.71
Mesonet 14.8 6.67 28.57 22.86 54.76 36.18 44.29 49.81 46.94 33.88
Xception 1.71 0 9.52 17.14 22.21 28.82 9.29 40.75 29.16 17.62
MultiAtten 4.13 0 28.57 13.81 44.42 30.32 35.96 43.93 43.57 27.19
CDCN++ 0.28 6.67 10.71 31.9 43.83 33.19 8.39 44.58 29.68 23.25
DeepPixel 0.14 0 10.12 13.33 40.19 32.8 2.14 37.38 19.48 17.29
ResNet50 1.21 0 6.55 12.86 28.79 21.55 4.29 36.82 23.06 15.01
ViT 0 0 0 9.52 46.3 16.2 2.68 30.99 20.01 13.97
JFSFDB-Uni 0 0 0.6 6.67 43.6 11.64 2.32 30.6 18.8 12.69
CLIP-V 0 0 10.71 8.33 37.17 23.46 11.43 25.72 27.81 16.8
CLIP 0 0 8.13 10.24 23.58 26.86 12.14 23.76 27.34 15.4
CoOp(Unified) 0.02 0 11.11 28.57 17.77 14.39 28.57 36.69 39.84 19.66
CoOp(Specific) 1.54 0 11.11 29.76 13.82 14.48 31.43 37.22 40.31 19.96
EER(%) Joint HipTune 0 0 0 3.84 11.72 8.99 1.23 24.92 12.21 6.99
TABLE X: Results on JFSFDB-EER.
JFSFDB primarily demonstrated that training with a combination of physical and digital attack samples can significantly enhance model performance. This conclusion is also applicable to our HiPTune model. When trained and tested separately within the dataset, HiPTune achieved an EER of 9.71% and an AUC of 93.12%. However, after joint training, HiPTune gained access to a more diverse set of classification criteria, which it could adaptively select and integrate, resulting in a substantial performance improvement: a 4.34% increase in AUC and a 2.72% reduction in EER. Moreover, the model’s generalization capability was significantly enhanced.
V-E Results on UniAttackData
Protocol Methods ACER ACC AUC EER Protocol Methods ACER ACC AUC EER
ResNet-50 1.35 98.83 99.79 1.18 ResNet-50 34.6 53.69 87.89 19.48
ViT-B/16 5.92 92.29 97 9.14 ViT-B/16 33.69 52.43 83.77 25.94
Auxiliary 1.13 98.68 99.82 1.23 Auxiliary 42.98 37.71 76.27 32.66
CDCN 1.4 98.57 99.52 1.42 CDCN 34.33 53.1 77.46 29.17
FFD 2.01 97.97 99.57 2.01 FFD 44.2 40.43 80.97 26.18
CLIP-V 0.46 99.55 99.95 0.46 CLIP-V 11.38 85.93 95.73 11.38
CLIP 0.58 97.4 99.93 0.58 CLIP 15.27 79.83 92.46 15.27
CoOp(unified) 7.94 90.96 97.9 7.94 CoOp(unified) 33.29 63.89 71.58 33.29
CoOp(specific) 8.49 90.98 97.45 8.49 CoOp(specific) 35.67 60.9 68.41 35.67
UniAttackDetection 0.52 99.45 99.96 0.53 UniAttackDetection 22.42 67.35 91.97 15.72
1 HiPTune 0.33 99.98 99.99 0.36 2 HiPTune 10.38 88.62 97.82 10.39
TABLE XI: Results on UniAttackData.
In the UniAttackData dataset, where the data distribution is relatively simple, our HiPTune model achieved outstanding performance. Leveraging the guidance function of prompts, HiPTune also delivered excellent results in cross-attack protocols, demonstrating its robustness and adaptability across different attack types.
V-F Abalation Study
To verify the effectiveness of the theory proposed by HiPTune and to analyze the influence of various settings, we conduct several ablation studies on protocol $P_{2}$ of HiFAB. These studies examine the depth of the tree, length of prompts, different selection strategies, various gate functions, and the contribution of each component. Additional experiments and detailed analyses of each element are provided in the Supplementary Materials.
Depth of Tree. The depth of the VP-Tree plays a crucial role in helping models understand fine-grained classification criteria, which enhances their ability to handle unseen advanced attacks during the integration process. Models such as CoOp, which lack a hierarchical structure, exhibit poor performance as they rely on a single, rigid classification criterion. In contrast, increasing the depth of the VP-Tree allows the model to ”draw” the line between live and fake faces adaptively and accurately, demonstrating the importance of a hierarchical approach.
Length of Prompts. The length of prompts determines the scope of learnable parameters, impacting both accuracy and efficiency. To explore this balance, we conduct several experiments to assess the effect of different prompt lengths. Using more parameters can capture more detailed classification-related information, enhancing model performance.
Component ACER(%) AUC(%) EER(%) ACC(%)
level-0 (CoOp) 14.20 92.57 14.26 85.44
level-1 9.78 94,26 10.12 90.12
level-2 5.78 96.88 5.45 95.08
level-3 3.26 98.78 3.67 95.79
TABLE XII: The influence of depth of the VP-Tree.
Component ACER(%) AUC(%) EER(%) ACC(%)
16 4.24 95.79 4.26 92.39
20 4.08 96.03 4.03 93.40
29 3.45 96.34 3.44 94.67
40 3.48 97.62 3.90 94.23
60 3.26 98.78 3.67 95.79
TABLE XIII: The impact on the length of prompts.
VI Conclusion
In this paper, we first present a large-scale Unified Face Attack Benchmark (HiFAB), which is the most extensive and comprehensive UAD dataset to date. It includes $14$ kinds of physical attacks, $40$ digital attacks, and $697,347$ videos in total. Every attack is applied to $2,875$ identities to reduce the ID noise. Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune), that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample. Finally, we conduct several experiments that evaluate the model’s performance in real-world scenarios.
VII Disscussion and Future Work
In this work, we explore the application of prompt tuning in the visual domain while acknowledging the value of examining its adaptability to its original text-based purpose. The hierarchical and semantic capabilities of text-based prompt tuning could provide valuable insights for visual applications, warranting further investigation. Furthermore, the suboptimal performance of baseline models under protocol $P_{1}$ may be partially attributed to the absence of a fine-tuning process on the testing datasets. This lack of adaptation limits the models’ ability to generalize to unseen attack types and domain variations. Further discussion on the impact of fine-tuning in bridging this gap will be presented.

### v3

@@ 5.2 Results on UniAttackData+
We employ a range of CLIP-based baselines, including CLIP [83], CLIP-V [83], and CoOp [86] to evaluate the performance in UAD on our dataset. In addition, we assess algorithms specifically designed for PAD/FFD, such as CFPL-FAS [90], MoE-FFD [91], and LLA-Net [92], to examine their generalization capabilities. We also compare our approach with the recently released FA3-CLIP, an algorithm tailored for unified attack detection. The detailed experimental results are presented in Tab. IV.
Protocol 1 is designed to evaluate the model’s generalization of face identity in the UAD task. As shown in Tab. IV, we can observe that CLIP-V [83] achieves decent performance, with an ACER of 9.36%. However, introducing fixed text templates or single-semantic learnable vectors does not bring performance benefits, such as the ACER of CLIP [83] and CoOp [86] being 9.93% and 14.76%, respectively. These experimental results indicate that in the UAD task, naive category prompts cannot accurately describe highly challenging multi-attack types, resulting in ineffective differentiation from real faces. Although the recently released CDCN++[89], CFPL-FAS[90], MoE-FFD [91], and LLA-Net [92] have achieved superior performance in PAD and FFD tasks, their effectiveness drops significantly on our UniAttackData+ dataset, with ACERs of 12.77%, 13.93%, 17.30%, and 10.57%, respectively. This suggests that these algorithms exhibit poor generalization across different attack categories, as models trained on PAD datasets fail to adapt effectively to FFD scenarios. FA3-CLIP [59] solves UAD by combining spatial and frequency features in a dual-stream architecture. By employing an attack-agnostic prompt learning and bridging the gap between different attack types, it significantly improves UAD performance, such as ACER is 8.26%. Our proposed HiPTune employs a visual prompt-based strategy that effectively guides the encoder toward informative features while preserving model efficiency. Furthermore, the hierarchical prompt tree captures multi-level, fine-grained classification criteria, enabling highly accurate and generalizable performance across diverse attack scenarios, with an ACER of just 7.31%.
Protocol 2 evaluates the generalization of the model across attack methods in the UAD task, which is widely applicable in real-world scenarios. When encountering unseen attacks, algorithms that either lack text prompts or rely solely on simple prompts exhibit significantly degraded performance. For example, CLIP-V [83] and CoOp [86] achieve ACERs of 28.77% and 43.14%, respectively. Although FA3-CLIP [59] demonstrates promising results on Protocol 1, its effectiveness significantly declines when encountering unknown attack types, with an ACER of 35.81%. This suggests that single-semantic prompt learning struggles to generalize across diverse attack categories. Unlike conventional prompt-matching methods, our prompts function solely as guidance rather than decision-making anchors, thereby reducing reliance on prior knowledge. This design enhances the model’s generalization to unseen attack methods while ensuring robustness and interpretability, achieving an ACER of 14.26%.
Protocol 3 is designed to evaluate whether models trained on simple attack types can retain high robustness when tested against complex and sophisticated attacks. We can observe that almost all methods have degraded performance, even to the point of random guessing, such as the ACER of CLIP-V [83], CoOp [86], MoE-FFD [91], and FA3-CLIP [59] exceeding 40%. In contrast, our HiPTune model adaptively integrates diverse types of prompts and demonstrates superior performance. Despite being trained on limited samples, HiPTune leverages high-level prompts and auxiliary information from other attack types to support inference. This design enables the model to maintain strong generalization capabilities, even against advanced attacks, achieving an ACER of 25.68%.
TABLE VI: Results on UniAttackData with standard Protocol 1, 1.1, 1.2, 1.3, and 2. ↓/↑ indicate that smaller/larger values correspond to better performance. The Avg. represents the mean ACER for Protocol 1, 1.1, 1.2, 1.3.
Methods P1 P1.1 P1.2 P1.3 Avg. P2
ACER(%)↓ AUC(%)↑ ACER AUC ACER AUC ACER AUC ACER AUC
ResNet-50 1.35 99.79 5.92 91.25 25.90 84.35 4.92 98.84 9.52 34.60 87.89
ViT-B/16 5.92 97.00 13.53 95.99 5.22 99.36 3.20 99.18 6.97 33.69 83.77
CDCN [89] 1.40 99.52 12.32 93.89 16.34 93.34 4.41 97.68 8.62 34.33 77.46
CLIP [83] 1.02 99.47 14.81 86.74 5.36 99.17 2.45 97.92 5.91 15.27 92.46
UniAttackDetection [39] 0.52 99.96 11.73 98.81 1.70 99.85 4.67 99.13 4.66 22.42 91.97
FA3-CLIP [59] 0.36 99.75 9.57 97.78 1.43 99.85 2.30 99.19 3.42 18.81 88.59
HiPTune 0.33 99.99 7.63 90.52 1.26 98.88 1.08 97.18 2.58 10.38 97.82
5.3 Results on JFSFDB Dataset
The Tab. V presents the evaluation results of various methods across three tasks: PAD, FFD, and UAD. Two metrics are used for performance comparison: Equal Error Rate (EER%) and Area Under Curve (AUC%). The experiment compares traditional methods (e.g., MesoNet [93], Xception [94], DeepPixel [6], CDCN++ [89], JFSFDB-Uni [38]), CLIP-based baselines (CLIP [83], CoOp [86], FA3-CLIP [59]), and the proposed method HiPTune.
In tasks PAD and FFD, our HiPTune achieves the best results in the AUC metric, such as 98.62%, 95.87%, and 98.62% on datasets MSU, 3DMask, and ROSE, and 89.72% and 89.63% on datasets DFDC and Celeb-DFv2, respectively. However, in terms of EER, HiPTune shows slightly weaker generalization on the ROSE and Celeb-DFv2 datasets. We attribute this to a class imbalance issue, where the model tends to learn better representations for the majority class (e.g., real faces). As a result, it achieves strong overall ranking performance (reflected in a high AUC). However, during testing, the decision threshold cannot effectively balance the trade-off between the two classes, leading to suboptimal EER performance. After training the UAD model with joint physics and digital forgery data, our HiPTune further improved the performance of AUC on most cross-dataset test results, such as MSU improving by 1.00% (98.62%→99.62%), DFDC improving by 1.16% (89.72%→90.88%), and Celeb-DFv2 improving by 3.72% (89.63%→93.35%). These experimental results indicate that our HiPTune can combine the diversity and complementarity of PAs and DFs to improve its performance in PAD and FFD tasks, respectively. However, for datasets 3DMask and ROSE, the performance of UAD slightly decreased, such as 95.87% decreasing to 95.43%, 98.62% decreasing to 97.82%, respectively. This indicates that introducing the FF++ dataset into the training set can impair the model’s ability to recognize 3D masks. We analyze that in our HiPTune framework, prompts are organized hierarchically by attack type. If a large influx of digital attacks (like from FF++) dominates the prompt learning phase, it could bias the learned prompts or pruning decisions, weakening the semantic alignment for physical branches (like 3D masks).
5.4 Results on UniAttackData
Tab. VI presents the performance of various methods on the UniAttackData under standard Protocols 1, 1.1, 1.2, 1.3, and 2, where lower ACER and higher AUC indicate better results. The Avg. column summarizes the average ACER across Protocols 1, 1.1, 1.2, and 1.3. HiPTune achieves the best overall performance, with the lowest average ACER (2.58%) and the highest AUC (99.99%) under Protocol 1, as well as strong performance in Protocol 2 (ACER: 10.38%, AUC: 97.82%). On Protocol 1, 1.2, and 1.3, HiPTune achieves the best ACER values of 0.33%, 1.26%, and 1.08%, respectively, and obtains the best mean ACER (2.58%). It indicates that HiPTune has the best generalization for facial identity information, such as the strict identity partitioning in the training, validation, and testing sets of Protocol 1.3. When physical presentation attacks and adversarial attacks are used as training sets, HiPTune has poorer generalization to deepfake attacks than ResNet-50, such as 7.63% v.s. 5.92% for ACER. On the contrary, the performance of ResNet-50 has significantly decreased, such as achieving an ACER value of 25.90% on Protocol 1.2. However, HiPTune still maintains its leading performance, achieving an ACER of 1.26%. This proves the effectiveness of using hierarchical prompt learning in HiPTune, where the model uses a hierarchical routing mechanism to select the most relevant prompts for each input image.
Finally, on the cross-attack category Protocol 2, where the model is trained on physical presentation attacks and tested on digital forgery (including opposite training and testing), our HiPTune demonstrates an absolute advantage, achieving an ACER of 10.38% and an AUC of 97.82%. HiPTune demonstrates strong generalization across different attack categories, due to the following points: (1) Hierarchical prompt modeling captures shared and distinct semantics. VP-Tree organizes attacks hierarchically by semantic similarity. High-level prompts (“physical” vs.“digital”) provide coarse but generalizable decision boundaries, while low-level prompts capture fine-grained modality-specific traits. This structure allows the model to transfer knowledge from one modality (e.g., physical) to another (e.g., digital), by reusing the shared higher-level decision boundaries. (2) Dynamic prompt selection tailored to each sample. The APP module ensures that, at inference time, the model can select the most appropriate semantic pathway, even if that pathway wasn’t explicitly optimized during training. This enables robust out-of-distribution generalization when tested on unseen attack types.
5.5 Results on cross-domain UAD benchmark
TABLE VII: Results on cross-domain UAD benchmark.
Method U&U+→J J&U+→U J&U→U+
ACER AUC ACER AUC ACER AUC
CLIP-V 45.12 56.26 13.15 91.86 39.81 63.4
CLIP 44.42 58.10 16.59 90.14 39.36 64.08
CoOp 38.21 65.84 30.28 76.11 29.77 76.77
HiPTune 32.93 70.89 8.55 94.67 25.93 83.05
Tab. VII presents the results of various methods, i.e., CLIP-V [83], CLIP [83], CoOp [86], and HiPTune on the cross-domain UAD benchmark, designed to evaluate the generalization ability of the models under domain shifts. The benchmark consists of three protocols, each involving training and testing across different combinations of datasets: UniAttackData (U), UniAttackData+ (U+), and JFSFDB (J). Among the compared methods, HiPTune achieves the best performance across all three cross-domain settings. For example, when the test set is J, U, and U+, our HiPTune achieves ACER values of 32.93%, 8.55%, and 25.93%, respectively, which are significantly better than the compared algorithms. The superiority of domain generalization in our analysis comes from the following two points: (1) HiPTune constructs a VP-Tree that models attack types in a coarse-to-fine hierarchy, enabling the model to distinguish diverse attack semantics across different domains. This structured representation ensures that even if an unseen domain presents a new variant of an attack, the model can still leverage higher-level prompts to make correct predictions. (2) The APP module dynamically selects the most appropriate prompts for each sample based on its visual features, rather than relying on fixed global parameters. This flexibility allows the model to adapt to distribution shifts without retraining, thereby improving generalization to unseen domains.
5.6 Ablation Studies
Effect of VP-Tree Depth. To evaluate the contribution of the hierarchical structure in VP-Tree, we conduct experiments on protocol 1 of UniAttackData+ by varying the depth of the tree from 1 to 3. As shown in Fig. 5, deeper trees consistently improve the performance across all metrics. Specifically, the full three-level VP-Tree achieves the best ACER (7.31%) and EER (7.61%), indicating that more fine-grained semantic grouping leads to more robust classification criteria. This demonstrates the benefit of modeling attack diversity hierarchically.
Effect of Prompt Length. We also study the impact of the prompt length in our framework. As presented in Fig. 5, increasing the length from 10 to 40 improves performance steadily. The model achieves optimal results at a length of 40, with an ACER of 7.31% and an AUC of 97.87%. Further increasing the length to 50 offers marginal gains in accuracy but slightly degrades ACER and EER, suggesting a trade-off between prompt expressiveness and overfitting. Thus, we adopt 40 as the default length in our experiments.
Figure 5: Ablation studies on the level of VP-Tree and the length of the prompt.
6 Conclusion
In this paper, we first present a large-scale Unified Face Attack Benchmark (HiFAB), which is the most extensive and comprehensive UAD dataset to date. It includes $14$ kinds of physical attacks, $40$ digital attacks, and $697,347$ videos in total. Every attack is applied to $2,875$ identities to reduce the ID noise. Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune) that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample. Finally, the proposed HiPTune demonstrates superior performance and generalization across three unified face forgery detection benchmarks.